# Detecting ionized bubbles around luminous sources during the reionization era using HI 21-cm signal

### v1

@@ 4.1 Foreground subtraction
where parameters $a$, $b$, $c$, and $d$ are constant. Then, we find the best fit polynomial along each line of sight corresponding a baseline and, subsequently, subtract them from the total simulated mock visibility signal.
Figure 6: This shows recovered visibilities after foreground subtraction for the uGMRT (top panels) and SKA1-Low (bottom panels) at same baselines and frequency considered in Figure 5.
Figure 6 presents the residual visibility at 175 MHz after the foreground contribution is subtracted from the total mock visibility for the uGMRT (top panels) and the SKA1-Low (bottom panels). The residual signal contains the HI 21-cm signal, the system noise, and the un-subtracted foreground, if any. To assess the reliability of foreground subtraction, we compare the combined input HI 21-cm signal and system noise with the extracted signal for two different baselines ${U=34\lambda}$ (top left panel) and ${230\lambda}$ (top right panel), as depicted in Figure 7. In the bottom panels of Figure 7, we show a similar comparison for SKA1-Low for two different baselines ${U=11\lambda}$ (bottom left panel) and ${U=254\lambda}$ (bottom right panel). The plots show that the residual visibility closely matches the input visibility data, with only minor deviations. These slight deviations are possibly due to some residual foreground contribution and subtraction of the HI 21-cm signal and the system noise during the foreground subtraction. Consequently, we can conclude that we have mostly subtracted the foreground data from the mock visibility data, making it ready for further analysis, such as matched filtering. We do not present similar results for the second scenario at $z=8.3$ as it is similar to the first scenario.
Figure 7: These plots compare the recovered visibility (red dashed lines) with the original visibility consists of the HI 21-cm signal and noise contributions (blue lines) for the uGMRT (top panels) and SKA1-Low (bottom panels) at same baselines and observing frequency considered in Figure 5. Here visibilities are plotted for every fourth frequency channel for visual clarity.
@@ 5 Results
We see that, although in most cases, the peak is in the range of $22-26$ Mpc for the uGMRT, some extreme realizations show peaks at higher/lower filter radii. The peak value of the SNR for the uGMRT also changes in the range of $\sim 3-7$. Due to much lower system noise and higher sensitivity, the SKA1-Low should be able to detect the ionized bubble with reduced observing time and higher SNR. The estimated bubble size is also highly accurate for the SKA1-Low, which we do not explicitly show here.
@@
The left panel of Figure 12 shows the mean SNR along with the $1-\sigma$ spread for $100$ noise realizations of the uGMRT, while the right panel shows the same for the SKA1-Low with $20$ realizations at $z=8.3$. Further, the peak location varies between $24$ and $36$ Mpc for the uGMRT for different realizations (Figure 11), which can be seen in Figure 13, that shows a histogram of peak locations for 100 realizations for the uGMRT. The peak value of the SNR for the uGMRT also changes in the range of $\sim 4-6$ and $6.5-8.5$ for the SKA1-Low.
@@ 5.4 Constraining the neutral hydrogen fraction
To calculate $C$, we simulate an entirely neutral and uniform HI 21-cm signal around a spherical ionized bubble of radius 23 Mpc, i.e., $x_{\rm HI}=1$ outside the ionized bubble. It allows us to estimate $C$ using eq. 5.1. Now, after adding the foreground component only with the HI 21-cm signal, we follow the entire analysis and calculate the SNR as a function of filter size without adding system noise from the radio interferometers. The SNR, in this case, peaks at 23 Mpc and is supposed to be equal to $C$ (following eq. 5.1) as the $x_{\rm HI}=1$ outside the ionized bubble in this case. We use this value of $C$ to estimate the mean $x_{\rm HI}$ and the $1\sigma$ uncertainty associated with it. The estimated neutral fraction is $x_{\rm HI}=0.93\pm 0.15$ at $z=7.1$ for $2048$ hrs of observations with uGMRT. It is very close to the actual $x_{\rm HI}$ in the simulation, which is $0.88$. The estimated $1\sigma$ upper limit of $x_{\rm HI}$ exceeds unity which is, of course, unrealistic. The SKA1-Low places much tighter constraints, which is $x_{\rm HI}=0.92\pm 0.055$ using only $96$ hrs of observations for the same scenario.
@@ 6 Summary & Discussion
First, we simulated the HI 21-cm signal around bright QSOs/galaxies matching the above cases and foreground contaminants in a 3D cube. Subsequently, we have generated visibilities corresponding to the HI 21-cm signal and foregrounds at baselines similar to the uGMRT and SKA1-Low experiments, and added the system noise with the visibilities corresponding to the HI signal and foreground contaminants. We then employ a foreground mitigation technique based on a polynomial fitting algorithm to remove them from the total visibilities. The residual visibilities match well with the combination of the input HI signal and system noise. However, it is heavily dominated by the system noise. To minimize the effects of system noise in the estimated quantity, we employed a technique based on the matched filter algorithm. This technique is applicable in our case as the functional form of the desired HI 21-cm signal around bright sources is known. We find that the expected signal-to-noise ratio for the proposed estimator for the uGMRT is $\sim 5$ for both $\sim 2000$ hours of observations at redshift $7.1$ and $\sim 3000$ hours of observations at redshift $8.3$. The SKA1-Low will have a much higher SNR ($\sim 8-10$) with just $\sim 100$ hrs of observations at same redshifts $7.1$ and $8.3$.
Apart from detecting the HI 21-cm signal, the technique is also able to measure the size of the ionized regions and HI neutral fraction. For the uGMRT $\sim 2000$ hrs of observations at redshift $7.1$, we find that the estimated bubble size ranges between $22-26$ Mpc for $100$ independent noise realizations when the actual bubble size is $23.5$ Mpc. Similar results are obtained for $\sim 3000$ hrs of observations with uGMRT at redshift $8.3$. The SKA1-Low should measure the bubble size more accurately, with the uncertainty in bubble size estimation being less than $1$ Mpc at both mentioned redshifts.
@@
The process of foreground subtraction using a polynomial along the frequency axis also removes some of the desired HI signal. We found that the fraction of subtracted HI signal increases when using smaller frequency bandwidth, significantly reducing the signal-to-noise ratio. We thoroughly investigated this effect using a wide range of bandwidths and found that the SNR increases by more than four times when the frequency bandwidth is increased from $5$ MHz to $16$ MHz for bubble size $23.5$ Mpc. We studied two particular cases here and results may vary if sizes of ionized bubbles and redshifts changes. However, our methods can be applied to any similar cases.
Here, we focused on the targeted detection of ionized bubbles around sources previously detected by other telescopes. Several such sources have been observed during the reionization epoch. Additionally, the matched filter technique can detect ionized bubbles even without prior source detection. However, this unguided detection requires exploring a vast parameter space, which is computationally expensive. The underlying bubble may not be spherical, as assumed in our study. However, as long as the bubble is relatively isolated, the method remains applicable, though the SNR might be reduced. Foregrounds are assumed to be spectrally smooth here, which may not be accurate in actual observations. A more sophisticated method for foreground mitigation may be needed in case foregrounds behave differently. We also ignored radio point sources, free-free etc. These practical issues will be addressed in future works.
@@ Acknowledgments
AM and CSM acknowledge financial support from Council of Scientific and Industrial Research (CSIR) via CSIR-SRF fellowships under grant no. 09/0096(13611)/2022-EMR-I and 09/1022(0080)/2019-EMR-I respectively. KKD acknowledges financial support from SERB (Govt. of India) through a project under MATRICS scheme (MTR/2021/000384). SM acknowledges financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Govt. of India through the Core Research Grant No. CRG/2021/004025. Authors acknowledge the use of computational resources at IUCAA, Pune.

### v2

@@ 4.1 Foreground subtraction
where parameters $a$, $b$, $c$, and $d$ are constant. Then, we find the best-fit polynomial along each line of sight corresponding to a baseline and subsequently subtract them from the total simulated mock visibility signal.
Figure 7: This shows recovered visibilities after foreground subtraction for the uGMRT (top panels) and SKA1-Low (bottom panels) at the same baselines and frequency considered in Figure 5.
Figure 7 presents the residual visibility at 175 MHz after the foreground contribution is subtracted from the total mock visibility for the uGMRT (top panels) and the SKA1-Low (bottom panels). Here, we have only considered the diffuse galactic synchrotron emission as the foreground contamination. However, we get a very similar residual signal when we add the radio point source contribution to the foreground contribution. The residual signal contains the HI 21-cm signal, the system noise, and the un-subtracted foreground, if any. To assess the reliability of foreground subtraction, we compare the combined input HI 21-cm signal and system noise with the extracted signal for two different baselines ${U=34\lambda}$ (top left panel) and ${230\lambda}$ (top right panel), as depicted in Figure 8. In the bottom panels of Figure 8, we show a similar comparison for SKA1-Low for two different baselines ${U=11\lambda}$ (bottom left panel) and ${U=254\lambda}$ (bottom right panel). The plots show that the residual visibility closely matches the input visibility data, with only minor deviations. These slight deviations are possibly due to some residual foreground contribution and subtraction of the HI 21-cm signal and the system noise during the foreground subtraction. Consequently, we can conclude that we have mostly subtracted the foreground data from the mock visibility data, making it ready for further analysis, such as matched filtering. We do not present similar results for the second scenario at $z=8.3$ as it is similar to the first scenario.
Figure 8: These plots compare the recovered visibility (red dashed lines) with the original visibility consisting of the HI 21-cm signal and noise contributions (blue lines) for the uGMRT (top panels) and SKA1-Low (bottom panels) at same baselines and observing frequency considered in Figure 5. Here visibilities are plotted for every fourth frequency channel for visual clarity.
@@ 5 Results
We see that, although in most cases, the peak is in the range of $22-26$ Mpc for the uGMRT, some extreme realizations show peaks at higher/lower filter radii. The peak value of the SNR for the uGMRT also changes in the range of $\sim 3-7$. Due to much lower system noise and higher sensitivity, the SKA1-Low should be able to detect the ionized bubble with reduced observing time and higher SNR. The estimated bubble size is also highly accurate for the SKA1-Low, which we do not explicitly show here. We note that the above analysis considers only the DGSE as the foreground contaminant. However, we have repeated the analysis including contributions from both the DGSE and radio point sources as foregrounds. We find very similar results to those presented here. Therefore, we do not explicitly show the outcomes of the combined foreground contamination here.
@@
The left panel of Figure 13 shows the mean SNR along with the $1-\sigma$ spread for $100$ noise realizations of the uGMRT, while the right panel shows the same for the SKA1-Low with $20$ realizations at $z=8.3$. Further, the peak location varies between $24$ and $36$ Mpc for the uGMRT for different realizations (Figure 12), which can be seen in Figure 14, which shows a histogram of peak locations for $100$ realizations for the uGMRT. The peak value of the SNR for the uGMRT also changes in the range of $\sim 4-6$ and $6.5-8.5$ for the SKA1-Low.
@@
5.3 Detectability of the QSO ionized bubble at $\bm{z=7.1}$ for $\bm{x_{\rm HI}=0.52}$
Here, we present our findings for the third case, where the targeted ionized bubble is surrounded by other ionized regions which result in a lower neutral hydrogen fraction of $\bm{x_{\rm HI}=0.52}$. Further, the targeted bubble at the center also became aspherical. A lower neutral hydrogen fraction means that the background HI 21-cm signal is weaker, which results in a lower SNR value. We use the same matched filtering technique as described earlier but increased the observation time to $\bm{5120}$ hours for the uGMRT and $\bm{200}$ hours for the SKA1-Low, considering the lower SNR and increased difficulty in detecting ionized regions at a lower neutral fraction. Figure 15 shows the SNR as a function of filter size $\bm{R_{\rm f}}$ for $\bm{5120}$ hours of uGMRT observation (left panel) and $\bm{200}$ hours of SKA1-Low observations (right panel) at $\bm{z=7.1}$. As in previous cases, each curve represents a different noise realization. We see that for the uGMRT, the SNR peaks at a filter size $\bm{\sim 25}$ Mpc, and for the SKA1-Low, it peaks at $\bm{\sim 23}$ Mpc, with minor deviations in different noise realizations.
Figure 15: This figure shows the signal-to-noise ratio for $\bm{10}$ independent noise realizations corresponding to $\bm{5120}$ hours of observation time for the uGMRT (left panel) and $\bm{200}$ hours of SKA1-Low observations (right panel) at redshift $\bm{7.1}$.
Further, we see that peak SNR values in all ten realizations are $\bm{\gtrsim 3}$ for the uGMRT and $\bm{\sim 11}$ for the SKA1-Low. Like the previous cases, we assume a total bandwidth of $\bm{16}$ MHz and $\bm{6.25}$ MHz while subtracting the foreground for the uGMRT and SKA1-Low, respectively. We see that it is possible to reliably detect individual ionized bubbles in the HI 21-cm maps using the uGMRT with $\bm{5120}$ hours of observation time and the SKA1-Low with only $\bm{\sim 200}$ hours of observations at $\bm{z=7.1}$ with significantly much lower neutral hydrogen fraction value.
Figure 16: The left panel shows the mean signal-to-noise (SNR) ratio of $\bm{100}$ independent noise realizations as a function of filter size for $\bm{5120}$ hours of the uGMRT observations at redshift $\bm{7.1}$. The right panel shows the SNR of $\bm{20}$ independent noise realizations for $\bm{200}$ hours of the SKA1-Low observations. The shaded region shows $\bm{1\sigma}$ uncertainty estimated from independent realizations.
[…]
The left panel of Figure 16 shows the mean SNR along with the $\bm{1-\sigma}$ spread for $\bm{100}$ noise realizations of the uGMRT, while the right panel shows the same for the SKA1-Low with $\bm{20}$ realizations at $\bm{z=7.1}$ for $\bm{x_{\rm HI}=0.52}$. Further, the peak location varies between $\bm{16}$ and $\bm{32}$ Mpc for the uGMRT for different realizations (Figure 15), which can be seen in Figure 17, that shows a histogram of peak locations for $\bm{100}$ realizations for the uGMRT. The peak value of the SNR for the uGMRT also changes in the range of $\bm{\sim 3-5}$ and $\bm{10.5-12.5}$ for the SKA1-Low.
@@ 5.5 Constraining the neutral hydrogen fraction
To calculate $C$, we simulate an entirely neutral and uniform HI 21-cm signal around a spherical ionized bubble of radius 23 Mpc, i.e., $x_{\rm HI}=1$ outside the ionized bubble. It allows us to estimate $C$ using eq. 5.1. Now, after adding the foreground component only with the HI 21-cm signal, we follow the entire analysis and calculate the SNR as a function of filter size without adding system noise from the radio interferometers. The SNR, in this case, peaks at 23 Mpc and is supposed to be equal to $C$ (following eq. 5.1) as the $x_{\rm HI}=1$ outside the ionized bubble in this case. We use this value of $C$ to estimate the mean $x_{\rm HI}$ and the $1\sigma$ uncertainty associated with it. The estimated neutral fraction is $x_{\rm HI}=0.93\pm 0.15$ at $z=7.1$ for $2048$ hrs of observations with uGMRT. It is very close to the actual $x_{\rm HI}$ in the simulation, which is $0.88$. The estimated $1\sigma$ upper limit of $x_{\rm HI}$ exceeds unity which is, of course, unrealistic. The SKA1-Low places much tighter constraints, which is $x_{\rm HI}=0.92\pm 0.055$ using only $96$ hrs of observations for the same scenario. Similarly we also studied this for the third case, where the neutral hydrogen fraction is lower in the simulation, discussed in subsection 5.3. We find the estimated neutral hydrogen fraction to be $\bm{x_{\rm HI}=0.49\pm 0.09}$ at $\bm{z=7.1}$ for $\bm{5120}$ hours of observations with the uGMRT, which is very close to the actual neutral hydrogen fraction $\bm{x_{\rm HI}=0.52}$.
@@ 6 Summary & Discussion
First, we simulated the HI 21-cm signal around bright QSOs/galaxies matching the above cases and foreground contaminants in a 3D cube. Subsequently, we have generated visibilities corresponding to the HI 21-cm signal and foregrounds at baselines similar to the uGMRT and SKA1-Low experiments and added the system noise with the visibilities corresponding to the HI signal and foreground contaminants. We then employ a foreground mitigation technique based on a polynomial fitting algorithm to remove them from the total visibilities. The residual visibilities match well with the combination of the input HI signal and system noise. However, it is heavily dominated by the system noise. To minimize the effects of system noise in the estimated quantity, we employed a technique based on the matched filter algorithm. This technique is applicable in our case as the functional form of the desired HI 21-cm signal around bright sources is known. We find that the expected signal-to-noise ratio for the proposed estimator for the uGMRT is $\bm{\sim 5}$ for both $\bm{\sim 2000}$ hours of observations at redshift $\bm{7.1}$ and $\bm{\sim 3000}$ hours at redshift $\bm{8.3}$, when the mean HI fraction outside the targeted bubble is $\bm{\sim 0.9}$. However, the required observing time increases to $\bm{\sim 5120}$ hrs for the uGMRT at redshift $\bm{7.1}$ to achieve a similar SNR when the mean HI fraction is lower at $\bm{\sim 0.5}$. The SKA1-Low will have a much higher SNR ($\sim 8-10$) with just $\sim 100$ hrs of observations at same redshifts $7.1$ and $8.3$, when the mean HI fraction outside the targeted bubble is $\bm{\sim 0.9}$. However, the required observing time increases to $\bm{\sim 200}$ hrs for the SKA1-Low at redshift $\bm{7.1}$ to achieve a similar SNR when the mean HI fraction is lower at $\bm{\sim 0.5}$.
Apart from detecting the HI 21-cm signal, the technique is also able to measure the size of the ionized regions and HI neutral fraction. For the uGMRT $\sim 2000$ hours of observations at redshift $7.1$, we find that the estimated bubble size ranges between $22-26$ Mpc for $100$ independent noise realizations when the actual bubble size is $23.5$ Mpc. Similar results are obtained for $\bm{\sim 5120}$ hours of observations with the uGMRT at redshift $\bm{7.1}$ with a mean HI fraction of $\bm{\sim 0.5}$ and for $\bm{\sim 3000}$ hours at redshift $\bm{8.3}$ with a mean HI fraction of $\bm{\sim 0.9}$. The SKA1-Low should measure the bubble size more accurately, with the uncertainty in bubble size estimation being less than $1$ Mpc at both mentioned redshifts.
@@
The process of foreground subtraction using a polynomial along the frequency axis also removes some of the desired HI signal. We found that the fraction of subtracted HI signal increases when using smaller frequency bandwidth, significantly reducing the signal-to-noise ratio. We thoroughly investigated this effect using a wide range of bandwidths and found that the SNR increases by more than four times when the frequency bandwidth is increased from $5$ MHz to $16$ MHz for bubble size $23.5$ Mpc. We studied three particular cases here and results may vary if sizes of ionized bubbles and redshifts change. However, our methods can be applied to any similar cases.
Here, we focused on the targeted detection of ionized bubbles around sources previously detected by other telescopes. Several such sources have been observed during the reionization epoch. Additionally, the matched filter technique can detect ionized bubbles even without prior source detection. However, this unguided detection requires exploring a vast parameter space, which is computationally expensive. The underlying bubble may not be spherical, as assumed in our study. However, as long as the bubble is relatively isolated, the method remains applicable, though the SNR might be reduced. Foregrounds are assumed to be spectrally smooth here, which may not be accurate in actual observations. A more sophisticated method for foreground mitigation may be needed in case foregrounds behave differently. We would also like to note that this study does not account for potential errors arising from improper calibration [64], ionospheric effects, or radio-frequency interference (RFI). These practical issues will be addressed in future works.
@@ Acknowledgments
AM and CSM acknowledge financial support from Council of Scientific and Industrial Research (CSIR) via CSIR-SRF fellowships under grant no. 09/0096(13611)/2022-EMR-I and 09/1022(0080)/2019-EMR-I respectively. KKD acknowledges financial support from SERB (Govt. of India) through a project under MATRICS scheme (MTR/2021/000384). SM acknowledges financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Govt. of India through the Core Research Grant No. CRG/2021/004025. The authors acknowledge the use of computational resources available to the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune, and to the Cosmology with Statistical Inference (CSI) research group at the Indian Institute of Technology Indore (IIT Indore).